OpenStudy (e.cociuba):

*****Hardest math problem in the world?*** Here it is: (9^62773 + 2)^83721 Yes thats (9 to the 62773 power plus 2) to the 83721 power. Now heres the rest of the problem...after finding the Huge number i have to find its digital sum. if u dont know what that means just give me the answer to the problem. Metal to the person who gets this!! ( i just cant find a calculator that can do this!!!)

OpenStudy (anonymous):

http://www.thonky.com/digital-root-calculator/

OpenStudy (anonymous):

I think the digital root is the same as the digital sum

OpenStudy (anonymous):

the answer is 3

OpenStudy (anonymous):

i did the problem took 2 hours but i did it

OpenStudy (anonymous):

Yeah.. that is what my link says

OpenStudy (e.cociuba):

i didnt even post this 2 hrs ago :)

OpenStudy (anonymous):

huh i answered it you just gave him a link

OpenStudy (anonymous):

How did it take you two hours if this question was posted 7 minutes ago? @desigordon

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

2 mins

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

Yeah, a link with the answer on it !!!! @desigordon

OpenStudy (koikkara):

itz logarithms..... (9^62773 + 2)^83721= 83721 x log (9^62773 + 2) log (9^62773 + 2) = log (9^62773) +log ( 2) log (9^62773 + 2) = 62773 log (9) + log ( 2) log (9^62773 + 2) = 62773 (0.954242509) + ( 0.301029996) log (9^62773 + 2) = 59900.966 So: 83721 x log (9^62773 + 2)= (83721)(59900.966) 83721 x log (9^62773 + 2) =5.01496877 × 10^9 Digital sum is 5 + 0+ 1+4+9+6+8+7+7= 47 4+7=11 1+1=2

OpenStudy (e.cociuba):

I got 2. Too! @Koikkara

OpenStudy (mertsj):

The log of the sum is NOT the sum of the logs!!!

OpenStudy (anonymous):

really so if i click the link and its blank the answer is trhere right

OpenStudy (anonymous):

Its 3 @Koikkara I think @e.cociuba

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

oh.. @desigordon It had it before

OpenStudy (anonymous):

i said the answer first so i win

OpenStudy (e.cociuba):

how is it 3??

OpenStudy (anonymous):

The digital root of (9^62773 + 2)^83721 is 3

OpenStudy (anonymous):

do you know logs

OpenStudy (anonymous):

The digital root of a number is the single digit that you get by adding all of the digits of the original number together. If the result of that is multiple digits, you add those digits together, repeating the process until you get a single digit. That digit is the digital root of the original number.

OpenStudy (e.cociuba):

k. yes i do

OpenStudy (e.cociuba):

k tnxs u guys!! Tht helped :)

OpenStudy (anonymous):

who gets medal

OpenStudy (e.cociuba):

I guess u do since u got the rite answer first!

OpenStudy (anonymous):

if you look at very top i said first answer

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

She, and your welcome

OpenStudy (anonymous):

i guess girls b4 guys so just give it to her she deserves it more than i do

OpenStudy (anonymous):

Thanks! ☺ @desigordon

OpenStudy (anonymous):

hey you guys mine helping me few a few problems

OpenStudy (e.cociuba):

Srry! @erdog82 my bad :) K then. Thts wat a man is all about @desigordon ! GBU :)

OpenStudy (e.cociuba):

Sure! Wtas up?

OpenStudy (anonymous):

how do you give medals

OpenStudy (e.cociuba):

U press 'Best Response' :)

OpenStudy (anonymous):

Solve the equation 3/2-1/x-4=-2/2x-8 . Use complete sentences to explain how you solved this equation.

OpenStudy (anonymous):

its a negative 2

OpenStudy (anonymous):

not positive

OpenStudy (mathstudent55):

@desigordon How do you know the digital root of (9^62773 + 2)^83721 is really 3? How did you calculate it?

OpenStudy (e.cociuba):

Ya! I'm not sure if the answer if really 3 @mathstudent55. Some help would be great!

OpenStudy (mathstudent55):

I have no idea what the answer is, but what I can tell you is that the link above www.thonky.com does not necessarily calculate it. It only looks at the digits you put in and ignores all other characters. If you enter a 2 in it, it returns 2. If you enter 23 it returns 5. So far, that's good because it's correct, but if you enter 2^3, it does not calculate that to be 8 and then return a value of 8. It will enter it as 23 (ignoring the ^ symbol) and will return a value of 5. Since the expression you are interested in has a sum and exponents, you are getting the root digit of the digits you are inputting and not of the enormous number the expression means.

OpenStudy (anonymous):

the real answer is 2